OpenStudy (anonymous):

Solve the equation: a.) 2 cos^2x-cos x=1

OpenStudy (anonymous):

let's define \[t:= \cos(x)\] we have the quadratic equation: \[2t^2-t=1\] can you find it's solutions?

OpenStudy (anonymous):

Would it be cos=1? Which means cos=0 degrees or 360 degrees?

OpenStudy (anonymous):

there are two solutions, \[\cos(x)=1\] and \[\cos(x)=-\frac{1}{2}\] so x can be either \[360^\circ{} k\] (for every integer k) or \[120^\circ{}+ 360^{\circ}k\] or \[240^\circ{}+ 360^{\circ}k\]